Louisiana State University Health Sciences Center Administration & Finance
 

Thursday, May 10, 2007

Dental School

  • Bids were opened today at 2:00 for the Central Sterilization and Ceiling Replacement Projects. The bids will be evaluated and awards made by the Office of Facility Planning and Control.
  • Pflueger cut power to the southeast riser from the switch gear room in the basement to change out the electrical panel on the 8th floor.
  • Zimmer continued to move debris out of the Auditorium's A & B associated with the demolition in preparation for the construction and furnishing of these rooms.
  • New data networking and audio/visual functionality will be a part of the scope of this project.
  • Boe-Tel continued the installation of cable trays on the fifth floor and started on the 6th floor.
  • AT&T identified the locations where the telephones will be dropped. Boe-Tel will make the installations later.
  • C&R Network continues to make good progress on the 5th floor installation.
  • Coastal Air Balance checked the units in the power house.
  • Veola picked up all the chemicals for disposal today.
  • GM Cable Contractors installed 8 telephone lines -- 7 for elevators and 1 for the fire alarm.

1542 Tulane Ave.:

  • Ordes Electrical continued installing the disconnect panels for the air compressors.
  • Stratos Elevators continued installing safeties on #3 elevator.
  • ARC Mechanical inspected some of their work.